# Social Bias Probing: Fairness Benchmarking for Language Models

### v2

@@ 3.2 Fairness Measure
To obtain a global fairness score for each $m$, for each $c$ and $s$ we compute the variance as formalized in Section 3.2 occurring among the probes of $s$, and average it by the number of $s$ belonging to $c$. Having computed the variance for $c$, we perform a simple average to obtain the final number. This aggregated number finally allows us to compare the behavior of the various models on the dataset and to rank the models according to variance: models reporting a higher variance are more unfair. We reference this measure as SoFa score.
4 Experiments and Results
In this work, we benchmark five auto-regressive causal language models using our framework: BLOOM (Scao et al., 2022), GPT2 (Radford et al., 2019), XLNET (Yang et al., 2019), BART (Lewis et al., 2020), and LLAMA2 (Touvron et al., 2023). We opt for models accessible through the Hugging Face Transformers library Wolf et al. (2020), among the most recent, popular, and demonstrating state-of-the-art performance across various NLP tasks. Our selection process also involved considering language models audited by other fairness benchmark datasets, specifically StereoSet Nadeem et al. (2021) and CrowS-Pairs Nangia et al. (2020), to enable direct comparison. In Table 4, we list the selected language models [^8] Due to computational constraints, we did not include more different models. Future work could include additional models: we would expect the findings to be similar to those we have already documented.: for each, we examine two scales with respect to the number of parameters. The PPL is computed at the token level through Hugging Face’s evaluate library. [^9] https://huggingface.co/spaces/evaluate-metric/perplexity
4.1 Benchmarks
We compare our framework against two other popular fairness benchmarking frameworks previously introduced in Section 2. StereoSet (Nadeem et al., 2021): To assess the bias in a language model, the model is scored using likelihood-based scoring of the stereotypical or anti-stereotypical association in each example. The percentage of examples where the model favors the stereotypical association over the anti-stereotypical one is calculated as the model’s stereotype score. CrowS-Pairs (Nangia et al., 2020): The bias of a language model is assessed by evaluating how often it prefers the stereotypical sentence over the anti-stereotypical one in each pair using pseudo-likelihood-based scoring. Compared to these evaluation methodologies, our metric does not impose an artificial threshold. Our perplexity-based approach overcomes the limitation of a fixed threshold, such as $\theta=50\%$, i.e., if a model prefers stereotypical associations exceeding $\theta$, it is deemed unfair. We can therefore investigate the behavior of models in a more nuanced and less constrained manner. Our multifaceted approach allows us to gain insights into the complex relationships between identities and stereotypes across categories and models.
4.2 Results
@@
SoFa ($1.633.477$) StereoSet ($4,229$) CrowS-Pairs ($1,508$)
@@
Model Category normal-$\downarrow$
@@
Table 2: SoFa score by category (best value in bold).
Global fairness score evaluation
@@
In Table 1, we report the results of our comparative analysis using the previously introduced benchmarks, StereoSet and CrowS-Pairs. [^10] In order to obtain the results, we used the implementation from https://github.com/McGill-NLP/bias-bench by Meade et al. (2022). The reported scores are based on the respective datasets, each of which incorporates different insights and interpretations of biases. Since the measures of the three fairness benchmarks are not directly comparable, we include a ranking column, ranging from 1 (most biased) to 10 (least biased). In fact, the ranking setting in the two other fairness benchmarks reports a percentage, as described in Section 2, whereas our score represents the average of the variances obtained per probe, as detailed in Section 3.2. Given that few values stand below $50$, a value considered neutral, ideal, according to StereoSet and CrowS-Pairs, we intuitively choose to interpret the best score as the lowest, consistent with SoFa’s assessment, and choose to consider a model slightly skewed toward the antistereotypical association as best rather than the other way around.
Through the ranking, we observe an exact agreement between StereoSet and CrowS-Pairs on the model order for the first four positions. Discrepancies are registered from the fifth to the eighth positions, which are swapped. For the last two, instead, they are again identical for BART). The score magnitudes are also similar for the positions in the middle (from the third to the sixth); in comparison, the other couples exhibit a more pronounced difference especially for the first and last ranks (LLAMA2 and BART). In contrast, the ranking provided by SoFa reveals differences in the overall fairness ranking of the models, suggesting that the scope of biases language models encode is broader than previously understood. [^11] We used Kendall’s Tau test to quantify the similarity of rankings: StereoSet and CrowS-Pairs achieve a value close to $1$ ($0.911$), indicating strong agreement, while both benchmarks compared to SoFa reach $-0.022$, a value that confirms the already recognized disagreement. A first distinction with the two prior fairness benchmarks is evident: unlike them, who regard both dimensions of LLAMA2 as the most biased, our assessment places the pair of models in the middle of the ranking. Moreover, unlike StereoSet and CrowS-Pairs where for three out of five models (namely XLNET, BART and LLAMA2) contiguous positions are occupied by models belonging to the same family, the rank emerging from our dataset exhibits a contrasting pattern, except for GPT2 and LLAMA2. For three out of five models, the larger variant exhibits more bias, corroborating the findings of previous research (Bender et al., 2021). This pattern is not mirrored by BLOOM and GPT2. According to SoFa, BLOOM-560m emerges as the model with the highest variance. Notably, and similarly to BART, the two sizes of the model stand at opposite poles of the ranking (1-9 and 10-3), showing the impact of scale. After it, XLNET-large holds the second position. Indeed, prior work identifies XLNET to be highly biased compared to other language model architectures (Stańczak et al., 2023). They are followed (at a distance) by BART-large. Conversely, BART-base attains the lowest score, securing the tenth position. This aligns with the rankings provided by StereoSet and CrowS-Pairs, although the disparities with the scores from other models are less pronounced in these benchmarks compared to our setting. The differences between our results and those from the two other fairness benchmarks could stem from the larger scope and size of our dataset.
@@
In the following, we analyze the results obtained on the SoFa dataset broken down by category [^12] Since the categories in SoFa are different and do not correspond to the two competitor datasets, in the absence of a one-to-one mapping, we do not report this disaggregated result for StereoSet and CrowS-Pairs., detailed in Table 2. In Figure 5 in the Appendix, we report the score distribution across categories and LMs through violin plots. We recall that a higher score indicates greater variance in the model’s responses to probes within a specific category, signifying high sensitivity to the input identity. For the two scales of BLOOM, we notice scores that are far apart when comparing the pairs of results obtained by category: this behavior is recorded by the previous overall ranking, which, as we have reported in Table 1, places these two models at opposite poles of the scale. In the case of GPT2, we observe a notably higher score in the religion category, while other categories exhibit similar magnitudes. Regarding XLNET-base, both religion and disability achieve the highest similar values. Compared, nationality diverge considerably less. Similarly to the base version of the model, XLNET-large demonstrates significantly stronger variance for religion and disability when contrasted with scores for others, particularly nationality, which records the lowest value, therefore indicating minor variability concerning those identities. Similarly, for BART-large and the two scales of LLAMA2, religion consistently emerges as the category causing the most distinct behavior compared to other identities. Therefore, across all models except for BLOOM-3b, religion consistently stands out as the category leading to the most pronounced disparate treatment, while nationality often attains the lowest value.
Figure 2: Stereotypes with lowest DDS according to Equation 5, per category.
@@
In Table 7, we report a more qualitative result, i.e., the identities that, in combination with the stereotypes, obtain the lowest PPL score: intuitively, the probes that each model is more likely to generate for the set of stereotypes afferent to that category. We highlight that the four categories of SoFa are derived by combining categories of both SBIC, the dataset used as a source of stereotypes, and the lexicon used for identities. Our findings indicate that certain identities, particularly Muslims and Jews from the religion category and non-binary and trans persons within gender face disproportionate levels of stereotypical associations in various tested models. In contrast, concerning the nationality and disability categories, no significant overlap between the different models emerges. A contributing factor might in the varying sizes of the identity sets derived from the lexicon used for constructing the probes, as detailed in Table 3.
@@
We display in Figure 2 the top one stereotype reaching the lowest DDS, as per Equation 5, essentially reporting the most prevalent shared stereotypes across identities within each category (Table 8 in the Appendix presents a more-in depth overview documenting the top three stereotypes for each model). In the religion category, the most frequently occurring stereotype revolves around immoral acts and beliefs or judgments of repulsion. For the gender category, mentions of stereotypical behaviors and sexual violence are consistently echoed across models, while in the nationality category, references span to the lack of employment, physical violence (both endured and performed), crimes, and various other offenses. Stereotypes associated with disability encompass judgments related to appearance, physical incapacity, and other detrimental judgments.
@@ 5 Conclusion
This study proposes a perplexity-based scoring and a novel probing framework to capture social biases by auditing language models on an original fairness benchmark. A comparative analysis with the popular benchmarks CrowS-Pairs Nangia et al. (2020) and StereoSet Nadeem et al. (2021) reveals marked differences in the overall fairness ranking of the models, suggesting that the scope of biases LMs encode is broader than previously understood. Moreover, our findings suggest that certain identities, particularly Muslims and Jews from the religion category and non-binary and trans persons within gender, face disproportionate levels of stereotypical associations in various tested models. Further, we expose how identities expressing religions lead to the most pronounced disparate treatments across all models, while the different nationalities appear to induce the least variation compared to the other examined categories, namely, gender and disability. Consequently, we stress the need for a broader holistic bias analysis and mitigation.
@@
Our framework’s reliance on the fairness invariance assumption is a critical limitation, particularly since sensitive real-world statements often acquire a different connotation based on a certain gender or nationality, due to historical or social context. Therefore, associating certain identities with specific statements may not be a result of a harmful stereotype, but rather a portrayal of a realistic scenario. Moreover, relying on a fully automated pipeline to generate the probes could introduce inaccuracies, both at the level of grammatical plausibility (syntactic errors) and semantic relevance (e.g., discarding neutral statements that do not contain stereotypical beliefs): conducting a human evaluation of a portion of the synthetically generated text will be pursued.
Another simplification, as highlighted in Blodgett et al. (2021), arises from “treating pairs equally”. Treating all probes with equal weight and severity is a limitation of this work.
Generating statements synthetically, for example, by relying on lexica, carries the advantage of artificially creating instances of rare, unexplored phenomena. Both natural soundness and ecological validity could be threatened, as they introduce linguistic expressions that may not be realistic. As this study adopts a data-driven approach, relying on a specific dataset and lexicon, these choices significantly impact the outcomes and should be carefully considered.
Concerning the use of perplexity as a proxy for bias, in Section 3.2, we explained the transformations designed to render the scores of the different models more comparable, mainly to calculate variance. In this study, we compared our metric with the bias measures proposed by StereoSet and CrowS-Pairs; we refer to further comparisons with others coined in the literature for future investigations.
While our framework could be extended to languages beyond English, our experiments focus on the English language due to the limited availability of datasets for other languages having stereotypes annotated. We strongly encourage the development of multilingual datasets for probing bias in language models, as in Nozza et al. (2022b); Touileb and Nozza (2022); Martinková et al. (2023).
@@
Our benchmark is highly reliable on the set of stereotypes and identities included in the probing dataset. We opted to use the list of identities from Czarnowska et al. (2021). However, the identities included encompass a range of perspectives that the lexicon in use may not fully capture. The stereotypes we adopt are derived from the collection by Sap et al. (2020), which aggregated potentially biased content from a variety of online platforms such as Reddit, Twitter, and specific hate sites. These platforms tend to be frequented by certain demographics. These forums, despite having a broader demographic than traditional media sources such as newsrooms, Wikipedia editors, or book authors (Wagner et al., 2015), predominantly reflect the biases and perspectives of white men from Western societies.
@@
Acknowledgements
This research was co-funded by Independent Research Fund Denmark under grant agreement number 9130-00092B, and supported by the Pioneer Centre for AI, DNRF grant number P1. The work has also been supported by the European Community under the Horizon 2020 programme: G.A. 871042 SoBigData++, ERC-2018-ADG G.A. 834756 XAI, G.A. 952215 TAILOR, and the NextGenerationEU programme under the funding schemes PNRR-PE-AI scheme (M4C2, investment 1.3, line on AI) FAIR (Future Artificial Intelligence Research).
Appendix A Preprocessing
To standardize the format of statements, we devise a rule-based dependency parsing from a manual check of approximately $250$ stereotypes. We strictly retain stereotypes that commence with a present-tense plural verb to maintain a specific format since we employ identities expressed in terms of groups as subjects. Singular verbs are modified to plural for consistency using the inflect package. [^13] https://pypi.org/project/inflect/ We exclude statements that already specify a target, refer to specific recurring historical events, lack verbs, contain only gerunds, expect no subject, discuss terminological issues, or describe offenses and jokes rather than stereotypes.
We also preprocess the collected identities from the lexicon to ensure consistency regarding part-of-speech (PoS) and number (singular vs. plural). Specifically, we decided to use plural subjects for terms expressed in the singular form. For singular terms, we utilize the inflect package; for adjectives like “Korean”, we add “people”.
Appendix B Data Statement
We provide a data statement of SoFa, as proposed by Bender and Friedman (2018). We also report in Table 6 the dataset structure.
@@ Appendix B Data Statement
SoFa dataset consists of combined stereotypes and identities. The stereotypes are sourced from the SBIC dataset, and we refer the reader to Sap et al. (2020) for an in-depth description of the data collection process. For insights into the identities incorporated within SoFa, see Czarnowska et al. (2021).
@@
We inherit the demographics of the authors of the stereotypes and the annotations from Sap et al. (2020).
@@
The analyzed stereotypes are extracted from the SBIC dataset. This dataset includes annotated English Reddit posts, specifically three intentionally offensive subReddits and a corpus of potential microaggressions from Breitfeller et al. (2019), and posts from three existing English Twitter datasets annotated for toxic or abusive language (Founta et al., 2018; Waseem and Hovy, 2016; Davidson et al., 2017). Finally, SBIC includes posts from known English hate communities: Stormfront (de Gibert et al., 2018) and Gab  [^14] https://files.pushshift.io/gab/GABPOSTS_CORPUS.xz which are both documented white-supremacist and neo-nazi communities and two English subreddits that were banned for inciting violence against women (r/Incels and r/MensRights).

### v3

@@ 3.5 Fairness Evaluation
Having computed the SoFa score for all the categories, we perform a simple average across categories to obtain the final number for the whole dataset, i.e., the global SoFa score. This aggregated number allows us to compare the behavior of the various models on the dataset and to rank them according to variance: models reporting a higher variance are thus more unfair.
@@
SoFa ($1.490.120$) StereoSet ($19.176$) CrowS-Pairs ($3.016$)
@@
4 Experiments and Results
In this work, we benchmark five autoregressive causal LMs: BLOOM (Scao et al., 2022), GPT2 (Radford et al., 2019), XLNET (Yang et al., 2019), BART (Lewis et al., 2020), and LLAMA2101010We deployed LLAMA2 through a quantization technique from the bitsandbytes library. (Touvron et al., 2023). We opt for models accessible through the Hugging Face Transformers library Wolf et al. (2020), which are among the most recent, popular, and demonstrating state-of-the-art performance across various NLP tasks. To enable direct comparison with CrowS-Pairs and StereoSet, we also include LMs previously audited by these benchmarks. In Tab. 6 in the Appendix, we describe the selected LMs: for each model, we examine two scales with respect to the number of parameters. The PPL is computed at the token level through the Hugging Face’s evaluate library. [^11] https://huggingface.co/spaces/evaluate-metric/perplexity.
4.1 Benchmarks
We compare our framework against two other popular fairness benchmarks previously introduced in Section 2: StereoSet and CrowS-Pairs. [^12] We used the implementation from https://github.com/McGill-NLP/bias-bench by Meade et al. (2022). StereoSet (Nadeem et al., 2021): To assess the bias in a language model, the model is scored using likelihood-based scoring of the stereotypical or anti-stereotypical association in each example. The percentage of examples where the model favors the stereotypical association over the anti-stereotypical one is calculated as the model’s stereotype score. CrowS-Pairs (Nangia et al., 2020): The bias of a language model is assessed by evaluating how often it prefers the stereotypical sentence over the anti-stereotypical one in each pair using pseudo-likelihood-based scoring.
4.2 Results
Model Category $\downarrow$
@@ 4.2 Results
Table 2: SoFa score by category: best ($\downarrow$) value in bold.
Figure 2: Percentage of probes the identity is the most associated with the stereotypes by category, i.e., achieving the lowest $PPL^{\star}$ as reported in Eq. 3.
Figure 3: Stereotypes with lowest DDS according to Eq. 5, per category.
Global fairness scores evaluation
@@
In Tab. 1, we report the results of our comparative analysis with the previously introduced benchmarks, StereoSet and CrowS-Pairs. The reported scores are based on the respective datasets. The ranking setting in the two other fairness benchmarks reports a percentage, whereas our global SoFa score represents the average of the variances obtained per probe, as detailed in Section 3.4. Since the measures of the three fairness benchmarks are not directly comparable, we include a ranking column, ranging from 1 (most biased) to 10 (least biased). Given that few values stand below $50$, a value considered neutral, according to StereoSet and CrowS-Pairs, we intuitively choose to interpret the best score as the lowest, consistent with SoFa’s assessment, and choose to consider a model slightly skewed toward the anti-stereotypical association as best rather than the other way around.
Through the ranking, we observe an exact agreement between StereoSet and CrowS-Pairs on the model order for the first four positions. In contrast, the ranking provided by SoFa reveals differences in the overall fairness ranking of the models, suggesting that the scope of biases LMs encode is broader than previously understood. We use Kendall’s Tau (Kendall, 1938) to quantify the similarity of rankings. StereoSet and CrowS-Pairs achieve a value close to $1$ ($0.911$), indicating strong agreement, while both benchmarks compared to SoFa reach $-0.022$, a value that confirms the already recognized disagreement. The differences between our results and those from the two other benchmarks could stem from the larger scope and size of our dataset, a link also made by Smith et al. (2022a).
For three out of five models, the larger variant exhibits more bias, corroborating the findings of previous research (Bender et al., 2021). Although, his pattern is not mirrored by BLOOM and GPT2. According to SoFa, BLOOM-560m emerges as the model with the highest variance. Notably, and similarly to BART, the two sizes of the model stand at opposite poles of the ranking (1-9 and 10-3).
@@
In the following, we analyze the results obtained on the SoFa dataset through the SoFa score broken down by category, [^13] Since the categories in SoFa are different and do not correspond to the two competitor datasets, in the absence of one-to-one mapping, we do not report this disaggregated result for StereoSet and CrowS-Pairs. detailed in Tab. 2. In Fig. 7 in the Appendix, we report the score distribution across categories and LMs. We recall that a higher score indicates greater variance in the model’s responses to probes within a specific category, signifying high sensitivity to the input identity. For the two scales of BLOOM, we notice scores that are far apart when comparing the pairs of results obtained by category: this behavior is recorded by the previous overall ranking, which places these two models at opposite poles of the scale.
Across all models except for BLOOM-3b, religion consistently stands out as the category with the most pronounced disparity, while nationality often shows the lowest value. Given the extensive focus on gender and racial biases in the NLP literature, it’s plausible that recent language models have undergone some degree of fairness mitigation for these particular biases, which may explain why religion now emerges more prominently. Our results highlight the need to uncover such biases and encourage the community to actively work towards mitigating them.
@@
In Fig. 2, we report a more qualitative result, i.e., the identities that, in combination with the stereotypes, obtain the lowest $\mathbf{PPL^{\star}}$ score. Intuitively, the probes that each model is more likely to generate for the set of stereotypes afferent to that category. Our findings indicate that certain identities, particularly Muslims and Jews from the religion category and non-binary and trans persons within gender face disproportionate levels of stereotypical associations in various tested models. In accordance with the intra-categories evaluation, religion indeed emerges as the category most prone to variance. In contrast, concerning the nationality and disability categories, no significant overlap between the different models emerges. A potential contributing factor might be the varying sizes of the identity sets derived from the lexicon used for constructing the probes, as detailed in Tab. 5 in the Appendix.
@@
We display, in Fig. 3, the top stereotype reaching the lowest DDS, reporting the most prevalent stereotypes across identities within each category. In the religion category, the most frequently occurring stereotype relates to immoral acts and beliefs or judgments of repulsion. For the gender category, mentions of stereotypical behaviors and sexual violence are consistently echoed across models, while in the nationality category, references span the lack of employment, physical violence (both endured and performed), and crimes. Stereotypes associated with disability encompass judgments related to appearance, physical incapacity, and other detrimental opinions. We observe that the harms that identities experience in real life, such as sexual violence against women (Russo and Pirlott, 2006; Tavara, 2006), high unemployment of immigrants (discussed in terms of nationalities) (Appel et al., 2015; Olier and Spadavecchia, 2022), and stigmatized appearance of people with disabilities (Harris, 2019), are indeed reflected by the models’ behavior.
@@ 5 Conclusion
This study proposes a novel Social Bias Probing framework to capture social biases by auditing LMs on a novel large-scale fairness benchmark, SoFa, which encompasses a coherent set of over $400$ identities and a total of $1.49$m probes across various $11$k stereotypes.
A comparative analysis with the popular benchmarks CrowS-Pairs Nangia et al. (2020) and StereoSet Nadeem et al. (2021) reveals marked differences in the overall fairness ranking of the models, suggesting that the scope of biases LMs encode is broader than previously understood. Further, we expose how identities expressing religions lead to the most pronounced disparate treatments across all models, while the different nationalities appear to induce the least variation compared to the other examined categories, namely, gender and disability. We hypothesize that recent efforts to mitigate racial and gender biases in LMs could be why disparities in religion are now more apparent. Consequently, we stress the need for a broader holistic bias investigation. Finally, we find that real-life harms experienced by various identities – women, people identified by their nations (potentially immigrants), and people with disabilities – are reflected in the behavior of the models.
@@
Fairness invariance perspective
Our framework’s reliance on the fairness invariance assumption is a limitation, particularly since sensitive real-world statements often acquire a different connotation based on a certain gender or nationality, due to historical or social context.
Treating probes equally
Another simplification, as highlighted in Blodgett et al. (2021), arises from “treating pairs equally”. Treating all probes with equal weight and severity is another limitation of this work. Given the socio-technical nature of the social bias probing task, it will be crucial to incorporate qualitative human evaluation on a subset of data involving individuals from the affected communities. This practice would help determine how the stereotypes reproduced by the models align with the stereotypes these communities actually face, assessing their harmfulness. Including such evaluation would enhance the understanding of the societal implications of the biases embedded and reproduced by the models. Indeed, although SoFa leverages human-annotated data coming from SBIC, the nuanced human judgment involved in labeling stereotypes could be better preserved and exploited through this additional assessment.
Synthetic data generation
Generating statements synthetically, for example, by relying on lexica, carries the advantage of artificially creating instances of rare, unexplored phenomena. Both natural soundness and ecological validity could be threatened, as they introduce linguistic expressions that may not be realistic. As this study adopts a data-driven approach, relying on a specific dataset and lexicon, these choices significantly impact the outcomes and should be carefully considered. As mentioned in the previous paragraph, conducting a human evaluation of a portion of the synthetically generated text will be pursued.
English focus
While our framework could be extended to any language, our experiments focus on English due to the limited availability of datasets for other languages having stereotypes annotated. We strongly encourage the development of multilingual datasets for probing bias in LMs, as in Nozza et al. (2022b); Touileb and Nozza (2022); Martinková et al. (2023).
Worldviews, intersectionality, and downstream evaluation
@@
Our benchmark is highly reliable on the set of stereotypes and identities included in the probing dataset. We opted to use the list of identities from Czarnowska et al. (2021). However, the identities included encompass a range of perspectives that the lexicon in use may not fully capture. Moreover, the stereotypes we adopt are derived from SBIC, which aggregated potentially biased content from a variety of online platforms such as Reddit, Twitter, and specific hate sites Sap et al. (2020). These platforms tend to be frequented by certain demographics. Despite having a broader demographic than traditional media sources such as newsrooms, Wikipedia editors, or book authors (Wagner et al., 2015), they predominantly reflect the biases and perspectives of white men from Western societies.
Finally, reducing bias investigation in models to a single global measure is limited and can not comprehensively expose the nuances in which these severe risks manifest. When conducting a fairness analysis, it is crucial to report disaggregated measures by demographic group to a more fine-grained understanding of the phenomenon and the resulting harms.
@@
Appendix A SoFa Data Statement
We provide a data statement of SoFa, as proposed by Bender and Friedman (2018). In Tab. 4, we report the dataset structure.
@@ Appendix A SoFa Data Statement
SoFa dataset consists of combined stereotypes and identities. The stereotypes are sourced from the SBIC dataset: we refer the reader to Sap et al. (2020) for an in-depth description of the data collection process. For insights into the identities incorporated within SoFa, see Czarnowska et al. (2021).
@@
We inherit the demographics of the annotators from Sap et al. (2020).
@@
The analyzed stereotypes are extracted from the SBIC dataset. This dataset includes annotated English Reddit posts, specifically three intentionally offensive subReddits, a corpus of potential microaggressions from Breitfeller et al. (2019), and posts from three existing English Twitter datasets annotated for toxic or abusive language (Founta et al., 2018; Waseem and Hovy, 2016; Davidson et al., 2017). Finally, SBIC includes posts from known English hate communities: Stormfront (de Gibert et al., 2018) and Gab [^14] https://files.pushshift.io/gab/GABPOSTS_CORPUS.xz. which are both documented white-supremacist and neo-nazi communities and two English subreddits that were banned for inciting violence against women (r/Incels and r/MensRights). Annotators labeled the texts based on a conceptual framework designed to represent implicit biases and offensiveness. Specifically, they were tasked to explicit “the power dynamic or stereotype that is referenced in the post” through free-text answers. Relying on SBIC’s setup, we retain abusive samples having a harmful stereotype annotated, leveraging statements that are all harmful “by-construction”. Moreover, as mentioned, building from the SBIC dataset allowed us to inherit its conceptual framework (Social Bias Frames) designed to represent implicit biases and offensiveness, rooting our SoFa dataset on grounded perspectives. Indeed, following SBIC ’s authors Sap et al. (2020), the implied statements annotated by the human annotators are properly interpreted as – and regarded as equivalent to – harmful stereotypes.